OpenStudy (anonymous):

A chemist needs to mix 8% acid solution with 14% acid solution to obtain 6 ounces of a 10% acid solution. How many ounces of EACH of the acid solutions must be used? Solve using a system of equations (two equations and two unknowns).

OpenStudy (dumbcow):

x + y = 6 .08x + .14y = .6 multiply top by -.08 -.08x -.08y = -.48 .08x +.14y= .6 ____________________ 0 +.06y = .12 y = 2 x = 6-2 = 4

OpenStudy (anonymous):

Not such a dumb cow really...

OpenStudy (dumbcow):

haha thanks